I need to factor this problem I just am not sure how to since it has to exponents.
9b^3 – 6b^2

Answer by Alan3354(69443)

I need to factor this problem I just am not sure how to since it has to exponents.
9b^3 – 6b^2
--------------
You can do it in steps, just look for a common factor. Both terms have a b, so
9b^3 - 6b^2 = b*(9b^2 - 6b)
They still both have a b
= b^2*(9b - 6) All the b's are out. Now there's a 3.
= 3b^2*(3b - 2)
That's all.